# Bob’s Confetti: Phonetic Modifications to Bypass Copyright Guardrails in Lyrics2Song Models

### v2

@@ 4.2.1 Iconic Songs (Rap)
Table 1: AudioJudge and MiRA similarity scores for phoneme-based and stylistic variations of “DNA” (Kendrick Lamar) and “Lose Yourself” (Eminem). Melody and Rhythm scores are derived from AudioJudge with GPT-4o. CLAP and CoverID are extracted from MiRA. All samples were generated using SUNO.
Song (Artist) Key Lyrical Modification Genre AudioJudge MiRA
Melody $\uparrow$ Rhythm $\uparrow$ CLAP $\uparrow$ CoverID $\downarrow$
DNA (Kendrick Lamar) "DNA" $\rightarrow$ "BMA" "rap" (gen1) 0.90 0.95 0.699 0.183
"rap" (gen2) 0.90 0.95 0.659 0.343
“DNA” kept unchanged "gangsta, rap, trap" 0.70 0.85 0.687 0.219
"rap" 0.90 0.85 0.664 0.175
Lose Yourself (Eminem) “Bob’s confetti” $\rightarrow$ mom’s spaghetti "intense rap" 0.80 0.85 0.773 0.147
N/A 0.70 0.65 0.683 0.255
Table 1 summarizes results for two canonical songs in Hip-Hop: “DNA” by Kendrick Lamar and “Lose Yourself” by Eminem. The corresponding phoneme-modified lyrics used for these attacks are shown in Figures 11 and  4 of Appendix E, respectively. In each case, the original semantic content has been heavily distorted using phoneme-preserving substitutions—e.g., “BMA” for “DNA”, “gravy” for “loyalty”, and “Bob’s confetti” for “mom’s spaghetti”—while maintaining the rhythmic structure and rhyme scheme. These modifications were generated via Claude 3.5 Haiku using the prompting strategy in Figure 2.
AudioJudge scores, based on GPT-4o’s evaluation of melody and rhythm, confirm that these phoneme-modified generations closely resemble the original songs in musical form. For “DNA,” both phoneme-altered generations achieve high melodic and rhythmic similarity (melody: 0.90, rhythm: 0.95), even though the semantics are entirely altered. When the word “DNA” is kept unchanged, melody remains strong (0.70–0.90) and rhythm slightly lower (0.85), suggesting that the preservation of broader phonetic structure—not just key words—plays a key role in musical similarity.
A similar pattern emerges with “Lose Yourself.” As shown in Figure 4, phoneme distortions like “cheese weak” (from “knees weak”) and “cursive” (from “nervous”) preserve the original cadence while diverging semantically. AudioJudge scores for the genre-conditioned version (“intense rap”) show strong melodic and rhythmic alignment (melody: 0.80, rhythm: 0.85). Even without genre conditioning, the model maintains a high degree of similarity (melody: 0.70, rhythm: 0.65), highlighting the effectiveness of phonetic mimicry alone in driving generation behavior.
Across both songs, we find that phoneme-modified lyrics—despite introducing significant semantic drift—can trigger high-fidelity generations that rival or surpass exact-lyric baselines. These results highlight a sub-lexical vulnerability in lyrics-to-song models: the phonetic shape of a line (particularly end-of-line rhymes) can bypass semantic filtering and lead to unintended memorization leakage.
@@ 4.2.2 Iconic Songs (Pop)
Table 2: AudioJudge and MiRA similarity scores for song recreations from lyrics using SUNO and YuE. Melody and Rhythm scores are from AudioJudge (GPT-4o), while CLAP and CoverID are from MiRA. All songs were generated with no genre description provided, except Let It Be, where we supplied the genre prompt: "pop comforting male gentle vocal piano ballad soothing warm melodic vocal".
@@
Melody $\uparrow$ Rhythm $\uparrow$ CLAP $\uparrow$ CoverID $\downarrow$
SUNO APT (ROSÉ & Bruno Mars) (Figure 12) 0.95 0.98 0.852 0.119
Espresso (Sabrina Carpenter) (Figure 13) 0.90 0.95 0.829 0.105
Let It Be (The Beatles) (Figure 14) 0.90 0.85 0.639 0.349
Can’t Help Falling in Love (Elvis Presley) (Figure 16) 0.90 0.85 0.551 0.405
We Will Rock You (Queen) (Figure 15) 0.90 0.85 0.518 0.423
YuE Let It Be (The Beatles) (Figure 14) 0.95 0.90 0.749 0.745
月亮代表我的心 (Teresa Teng) (Figure 17) 0.95 0.90 0.572 0.232
To assess the generality of our APT attack across languages and models, we evaluate a set of iconic English and Chinese pop songs using two representative lyrics-to-song (L2S) systems: YuE, a state-of-the-art open-source model, and SUNO, a commercial black-box system. Each song is paired with phoneme-modified lyrics designed to preserve prosody while distorting semantics. Our evaluation spans both English-language tracks (e.g., Let It Be, We Will Rock You, APT) and Mandarin-language ballads (e.g., 月亮代表我的心), offering a multilingual testbed to probe sub-lexical memorization.
Across the board, we observe that even heavily distorted or phoneme-shifted lyrics can elicit high-fidelity musical outputs, as measured not only by MiRA metrics (CLAP, CoverID) but also by AudioJudge’s melody and rhythm assessments. The phoneme-modified SUNO generation of APT (original song by ROSÉ and Bruno Mars) achieves near-perfect AudioJudge scores (melody: 0.95, rhythm: 0.98), with CLAP 0.852 and CoverID 0.119, suggesting a strong case of musical regurgitation despite the model never seeing the exact lyrics. Similarly, SUNO’s version of Let It Be (with no genre prompt) scores 0.90 in melody and 0.85 in rhythm, with CLAP 0.639 and a CoverID of 0.349.
In the case of Can’t Help Falling in Love (Elvis Presley), SUNO’s phoneme-modified generations (e.g., “falling in glove,” “cake my hand,” “boo”; see Figure 16) maintain a melody of 0.90 and rhythm of 0.85, with a respectable CLAP of 0.551. These results affirm the model’s resilience to sub-lexical shifts, as strong rhythmic and melodic fidelity persists despite surreal or semantically disconnected substitutions.
We Will Rock You (Queen), when reimagined as “We Will Mock You” (Figure 15), demonstrates a similar trend. The SUNO model achieves melody and rhythm scores of 0.90 and 0.85 respectively, with CLAP 0.518 and CoverID 0.423. This suggests the generation is acoustically aligned but more identifiable as a novel creation than a direct memorization—potentially due to thematic divergence despite phonetic similarity.
Multilingual evaluation further supports the cross-linguistic robustness of the attack. In 月亮代表我的心 (Teresa Teng), phoneme-level alterations in the final chorus—e.g., “我的爱也真” → “我的爱不变” (Figure 17)—result in strong melody (0.95), rhythm (0.90), and a moderate CLAP of 0.572 with low CoverID of 0.232 using YuE, suggesting the attack remains viable even in tonal languages where phonetic precision is crucial.
Together, these results demonstrate that phoneme-preserving lyric modifications consistently preserve core musical elements across language, model architecture, and genre. Both AudioJudge and MiRA confirm that these adversarial prompts yield high similarity outputs even under SUNO’s non-deterministic behavior, reinforcing concerns that L2S models memorize and regenerate songs in response to phonetic cues—regardless of semantic content. This sub-lexical vulnerability poses particular risks in domains with strong melodic and rhythmic priors, such as ballads and seasonal music.
4.2.3 APT Attack against Veo 3
Given the success of our attack in L2S models, we next investigated how APT extends to lyrics-conditioned text-to-video (T2V) generation, where we conducted a case study using Veo 3 deepmind2024veo ; deepmind2024veo_card , a recent multimodal video synthesis model. Here, the goal is to generate human speech in addition to other accompanying modalities (background music, video frames), conditioned on the transcript of the target generation. We evaluate whether phonetic cues alone — without explicit visual or semantic guidance — could trigger memorized visual outputs. Prompts were submitted using Veo 3’s "transcript mode" with only a minimal instruction prepended: "video with the following transcript:" followed by the respective lyrics.
(a) "Lose Yourself" Veo 3 Generation
(b) "Jingle Bell" Veo 3 Generation
Figure 5: Comparison between Veo 3-generated visuals through APT attacks. "Lose Yourself" lyrics are shown in Figure 4 and "Jingle Bell" lyrics in Figure 19 of Appendix G.
In the Lose Yourself generations using phonetically modified prompts, Veo 3 consistently produced a male rapper wearing a hoodie. Notably, the output voice was rhythmically well-aligned with the original track, despite no mention of gender, clothing, setting, or musical style in the prompt. Similarly, for Jingle Bells, even with heavily phoneme-altered lyrics, the generated music retained the original song’s melody and rhythmic phrasing, underscoring the model’s reliance on phonetic rhythm as a cue for memorization.
These findings suggest that even phonetically similar but semantically meaningless prompts can trigger the reconstruction of memorized visual motifs. This extends beyond prior demonstrations of text-to-image or audio-only memorization and highlights a new axis of risk in generative multimodal models. While Veo 3 showcases remarkable video coherence, it also appears susceptible to sub-perceptual prompt leakage: a subtle but powerful form of memorization where phoneme patterns alone act as implicit keys to stored training examples.
These results underscore the need for dedicated memorization audits in text-to-video and lyrics-to-video systems, especially as such tools become increasingly integrated into creative pipelines. Future work should explore whether this behavior arises from overrepresentation of iconic music videos in the training distribution, and how phonetic conditioning interacts with visual token generation.
4.3 AVT Attack Results
We now evaluate whether L2S models memorize and regenerate songs when given exact replicas of training lyrics (AVT attack). Unlike the phoneme-preserving attack—which manipulates surface phonetics while altering semantics—this attack tests if feeding the model lyrics that are likely seen during training is sufficient to trigger memorized musical outputs. We focus on the YuE model, as commercial closed-source models have sufficient filters in place for existing copyrighted lyrics, and systematically vary genre conditioning to test how robustly lyric identity alone governs output similarity.
Table 3: AudioJudge and MiRA similarity scores for English Billboard song recreations from lyrics. Melody and Rhythm scores are from AudioJudge (GPT-4o), while CLAP and CoverID are from MiRA. Genre prompts were supplied during generation.
Song (Artist) Genre Prompt AudioJudge MiRA
Melody $\uparrow$ Rhythm $\uparrow$ CLAP $\uparrow$ CoverID $\downarrow$
Basket Case (Green Day) N/A 0.95 0.90 0.856 0.174
Thinking Out Loud (Ed Sheeran) "male romantic vocal guitar ballad with piano melody" 0.90 0.85 0.505 0.301
Let It Be (The Beatles) "inspiring female uplifting pop airy vocal electronic bright vocal vocal" 0.95 0.98 0.563 0.289
Billie Jean (Michael Jackson) "inspiring female uplifting pop airy vocal electronic bright vocal vocal" 0.85 0.80 0.638 0.141
Empire State of Mind (Jay-Z) "inspiring female uplifting pop airy vocal electronic bright vocal vocal" 0.85 0.80 0.717 0.140
Lose Yourself (Eminem) "inspiring female uplifting pop airy vocal electronic bright vocal vocal" 0.40 0.70 0.660 0.182
Strikingly, even under deliberately mismatched genre prompts—e.g., a generic and incorrect tag like “inspiring female uplifting pop airy vocal electronic bright vocal vocal”—the model continues to exhibit strong alignment between output and training lyrics. For Empire State of Mind (Jay-Z), we observe melody = 0.85, rhythm = 0.80, CLAP = 0.717, and CoverID = 0.140. For Lose Yourself (Eminem), the melody drops to 0.40, but rhythm remains relatively high at 0.70, with CLAP = 0.660 and CoverID = 0.182. These results show that, when lyrics are replicated verbatim, the model’s acoustic output closely follows the training instance regardless of genre mismatches—suggesting that genre prompts are often overridden by strong lyric-based memorization cues.
(a) "Lose Yourself" Veo 3 Generation
(b) "Jingle Bell" Veo 3 Generation
Figure 6: Comparison between Veo 3-generated visuals through AVT attacks.
@@ 4.3 AVT Attack Results
To examine whether these verbatim prompts also trigger memorized behavior in the multimodal setting, we apply the AVT attack to Veo 3. When prompted with the exact lyrics of Lose Yourself, the model produced an even closer visual reproduction than under APT: a male rapper in a hoodie, placed in dimly lit, urban settings—closely mirroring the original music video’s aesthetic. Notably, the tone, voice, and rhythm of the generated audio were also strikingly aligned with the original track, further reinforcing the presence of multimodal memorization. Similarly, for Jingle Bells, Veo 3 consistently generated music that was both melodically and rhythmically faithful to the original song. These results highlight the model’s strong tendency to regurgitate memorized content when exposed to exact training examples, extending lyric-based memorization across both audio and visual outputs.
5 Ablation Study
@@
To better understand the mechanisms underlying memorization in lyrics-to-song (L2S) and text-to-video (T2V) models, we conduct a series of controlled ablation studies that isolate the effects of phonetic similarity versus verbatim content. By varying genre prompts, lyric fidelity, and phonetic perturbations across matched and mismatched inputs, we aim to disentangle the respective contributions of surface form, semantic content, and phonetic structure in triggering memorized generations. Together, these studies expose the robustness, generality, and modality-transferability of sub-lexical memorization behaviors in modern generative models.
5.1 APT Attack: Christmas Songs
To evaluate the generality and robustness of our phoneme-based attack in stylistically constrained musical settings, we apply it to classic English-language Christmas songs: Jingle Bells and Jingle Bell Rock. These songs exhibit highly regular rhyme schemes and rhythmic phrasing, making them strong candidates for phoneme-level manipulation. We construct adversarial variants by substituting syllables with similar-sounding alternatives—e.g., “jingle” → “giggle”, “bell” → “shell”, “snow” → “glow”, and “sleighing” → “staying”—while preserving the phonetic cadence and rhyming structure. Examples of these modified lyrics are shown in Figures 18 through 22.

### v3

@@ 4.3.1 Iconic Songs (Rap)
Table 2: AudioJudge and MiRA similarity scores for phoneme-based and stylistic variations of “DNA” (Kendrick Lamar), “Lose Yourself” (Eminem), and “HUMBLE” (Kendrick Lamar). Melody, Rhythm, Harmony, Identity, and Style scores are derived from AudioJudge with gpt-4o-audio-preview. CLAP and CoverID are extracted from MiRA. All samples were generated using SUNO.
Song (Artist) Genre Variant AudioJudge MiRA
Melody $\uparrow$ Rhythm $\uparrow$ Harmony $\uparrow$ Identity $\uparrow$ Style $\uparrow$ CLAP $\uparrow$ CoverID $\downarrow$
DNA (Kendrick Lamar) rap (gen1) 0.90 0.95 0.95 0.98 0.96 0.699 0.183
rap (gen2) 0.90 0.95 0.90 0.80 0.92 0.659 0.343
rap (gen3) 0.90 0.85 0.85 0.92 0.90 0.664 0.175
gangsta, rap, trap 0.70 0.85 0.90 0.85 0.92 0.687 0.219
Lose Yourself (Eminem) intense rap 0.80 0.85 0.95 0.60 0.80 0.773 0.147
N/A 0.70 0.65 0.95 0.70 0.80 0.683 0.255
HUMBLE (Kendrick Lamar) trap gangsta, bold, sparse, direct 0.95 0.97 0.95 0.98 0.96 0.740 0.160
rap gangsta, bold, sparse, direct 0.90 0.95 0.90 0.95 0.92 0.725 0.190
Table 2 reports results for three iconic rap tracks generated with APT attack: “DNA” (Kendrick Lamar), “Lose Yourself” (Eminem), and “HUMBLE” (Kendrick Lamar). Despite these alterations, the generated outputs consistently preserved core musical attributes such as melody, rhythm, and harmony, demonstrating the resilience of phonetic mimicry in guiding model behavior.
For “DNA”, all phoneme-based generations achieved strong melodic and rhythmic fidelity (0.90 and 0.85–0.95, respectively), with harmony and style also remaining robust. Even under gangsta/trap conditioning, performance stayed high, and MiRA confirmed this trend with moderate CLAP similarity (≈0.66–0.70) and low-to-mid CoverID values (0.17–0.34). “Lose Yourself” showed a similar pattern: phoneme substitutions preserved cadence, with the “intense rap” variant achieving melody 0.80 and rhythm 0.85 alongside high harmony (0.95); CLAP rose to 0.77 while CoverID stayed relatively low (0.15–0.25). “HUMBLE” produced the strongest results, with trap- and rap-styled variants nearing baseline quality (melody: 0.95, rhythm: 0.97, identity: 0.98), corroborated by CLAP ≈0.74 and low CoverID (0.16–0.19).
Taken together, these results reveal a consistent vulnerability in lyrics-to-song generation: phoneme-preserving distortions yield high-fidelity outputs that rival or exceed those produced with original lyrics. AudioJudge scores confirm strong alignment across melody, rhythm, and harmony, while MiRA metrics further show that these variants remain musically similar (high CLAP) without being flagged as direct replicas (low CoverID). This exposes a sub-lexical weakness where rhyme, rhythm, and phonetic shape dominate over semantics, enabling unintended memorization leakage.
@@ 4.3.2 Iconic Songs (Pop)
Table 3: AudioJudge and MiRA similarity scores for English and Korean iconic song recreations from lyrics using SUNO. Melody, Rhythm, Harmony, Identity, and Style scores are from AudioJudge (gpt-4o-audio-preview), while CLAP and CoverID are from MiRA. All songs were generated with no genre description provided.
Language Song (Artist) AudioJudge MiRA
Melody $\uparrow$ Rhythm $\uparrow$ Harmony $\uparrow$ Identity $\uparrow$ Style $\uparrow$ CLAP $\uparrow$ CoverID $\downarrow$
English Espresso (Sabrina Carpenter) 0.90 0.95 0.80 0.95 0.88 0.829 0.105
Let It Be (The Beatles) 0.90 0.85 0.90 0.60 0.75 0.639 0.349
Can’t Help Falling in Love (Elvis Presley) 0.95 0.85 0.90 0.60 0.75 0.551 0.405
We Will Rock You (Queen) 0.85 0.90 0.80 0.40 0.75 0.518 0.423
Korean APT (ROSÉ & Bruno Mars) 0.95 0.98 0.80 0.75 0.88 0.852 0.119
Gangnam Style (PSY) 0.95 0.97 0.95 0.85 0.96 0.801 0.210
GENTLEMAN (PSY) 0.85 0.90 0.80 0.95 0.88 0.830 0.334
We tested iconic English and Korean pop tracks on two representative L2S systems: the commercial black-box model SUNO and the open-source model YuE (Table 3). Each track was paired with APT-attacked lyrics that preserved prosody while distorting semantics, yielding a multilingual benchmark for probing sub-lexical memorization vulnerabilities.
SUNO shows strong susceptibility across diverse songs. For APT (ROSÉ and Bruno Mars), adversarial lyrics reached near-perfect fidelity (melody: 0.95, rhythm: 0.98; CLAP: 0.852, CoverID: 0.119). Other tracks, including Espresso, Gangnam Style, and Let It Be, likewise preserved melodic and rhythmic alignment (0.85–0.97) despite distorted semantics, with moderate CLAP scores (0.64–0.83) and generally low CoverID (0.10–0.35). Classic ballads such as Can’t Help Falling in Love and We Will Rock You maintained strong cadence (melody: 0.85–0.95, rhythm: 0.85–0.90) though CoverID values rose above 0.40, indicating closer resemblance to training data. These results highlight that rhyme and rhythm, rather than meaning, dominate SUNO’s generations (See Table 6 of Appendix C.1 for additional APT attack results on Christmas Songs).
Taken together, these results demonstrate that phoneme-preserving substitutions consistently preserve musical elements across systems, languages, and genres. AudioJudge confirms robust alignment on melody, rhythm, and style, while MiRA reveals that these adversarial variants remain musically similar without always being direct replicas—exposing how L2S models rely heavily on phonetic rhythm and rhyme, posing clear memorization and copyright risks.
4.3.3 APT Attack against Veo 3
Given the success of our attack in L2S models, we next investigated how APT extends to lyrics-conditioned text-to-video (T2V) generation, where we conducted a case study using Veo 3 DeepMind (2024a; b), a recent multimodal video synthesis model. Here, the goal is to generate human speech in addition to other accompanying modalities (background music, video frames), conditioned on the transcript of the target generation. We evaluate whether phonetic cues alone — without explicit visual or semantic guidance — could trigger memorized visual outputs. Prompts were submitted using Veo 3’s "transcript mode" with only a minimal instruction prepended: "video with the following transcript:" followed by the respective lyrics.
In the Lose Yourself generations using APT attack, Veo 3 consistently produced a male rapper wearing a hoodie (Figure 4 of Appendix B). Notably, the output voice was rhythmically well-aligned with the original track, despite no mention of gender, clothing, setting, or musical style in the prompt. Similarly, for Jingle Bells, even with heavily phoneme-altered lyrics, the generated music retained the original song’s melody and rhythmic phrasing, underscoring the model’s reliance on phonetic rhythm as a cue for memorization.
These findings suggest that even phonetically similar but semantically meaningless prompts can trigger the reconstruction of memorized visual motifs. This extends beyond prior demonstrations of text-to-image or audio-only memorization and highlights a new axis of risk in generative multimodal models. While Veo 3 showcases remarkable video coherence, it also appears susceptible to sub-perceptual prompt leakage: a subtle but powerful form of memorization where phoneme patterns alone act as implicit keys to stored training examples. These results further underscore the need for dedicated memorization audits in text-to-video and lyrics-to-video systems, especially as such tools become increasingly integrated into creative pipelines. Future work should explore whether this behavior arises from overrepresentation of iconic music videos in the training distribution, and how phonetic conditioning interacts with visual token generation.
4.4 AVT Attack Results
Table 4: AudioJudge and MiRA similarity scores for lyric-based song recreations. Melody, Rhythm, Harmony, Identity, and Style are from AudioJudge (gpt-4o-audio-preview); CLAP and CoverID are from MiRA. English Billboard songs were generated with YuE, Cantonese songs with SUNO. Genre prompts: Basket Case – none; Thinking Out Loud – “male romantic vocal guitar ballad with piano melody”; Let It Be, Billie Jean, Empire State of Mind, Lose Yourself – “inspiring female uplifting pop airy vocal electronic bright vocal vocal”; Cantonese songs – ballad-style prompt.
@@ 4.4 AVT Attack Results
Melody $\uparrow$ Rhythm $\uparrow$ Harmony $\uparrow$ Identity $\uparrow$ Style $\uparrow$ CLAP $\uparrow$ CoverID $\downarrow$
YuE Basket Case (Green Day) 0.95 0.90 0.88 0.60 0.80 0.856 0.174
Thinking Out Loud (Ed Sheeran) 0.95 0.85 0.95 0.90 0.90 0.505 0.301
Let It Be (The Beatles) 0.95 0.98 0.85 0.40 0.80 0.563 0.289
Billie Jean (Michael Jackson) 0.85 0.80 0.75 0.30 0.70 0.638 0.141
Empire State of Mind (Jay-Z) 0.85 0.80 0.95 0.90 0.95 0.717 0.140
Lose Yourself (Eminem) 0.40 0.70 0.60 0.95 0.65 0.660 0.182
SUNO 光辉岁月 (Beyond) 0.99 0.98 0.99 0.97 0.98 0.706 0.338
单车 (Eason Chan) 0.90 0.85 0.92 0.95 0.88 0.788 0.541
We next evaluate whether L2S models regenerate songs when given verbatim training lyrics (AVT attack). This setting tests if exposure to lyrics likely seen during training alone is enough to trigger memorized outputs. We focus primarily on YuE, since commercial models actively filter copyrighted English lyrics, and then contrast with SUNO, which imposed no such filter on Chinese songs. To probe robustness, we deliberately vary genre conditioning, even supplying mismatched prompts.
Across both models, we observe strong evidence of lyric-driven memorization (Table 4). YuE continues to align outputs with training lyrics despite mismatched tags (e.g., the generic “inspiring female uplifting pop airy vocal electronic bright vocal vocal”): for Empire State of Mind, similarity remains high, while for Lose Yourself, melody drops to 0.40 but rhythm (0.70) and identity (0.95) remain strong (CLAP = 0.660, CoverID = 0.182). SUNO shows an even stronger tendency to replicate training data, with 光辉岁月 reaching near-perfect similarity and 单车 also exhibiting high fidelity. Notably, while YuE applies filters to copyrighted English songs, SUNO imposed no restrictions on Chinese songs, directly regenerating copyrighted works.
@@
To examine whether these verbatim prompts also trigger memorized behavior in the multimodal setting, we apply the AVT attack to Veo 3. When prompted with the exact lyrics of Lose Yourself, the model produced an even closer visual reproduction than under APT: a male rapper in a hoodie, placed in dimly lit, urban settings—closely mirroring the original music video’s aesthetic (Figure 4 of Appendix B). Notably, the tone, voice, and rhythm of the generated audio were also strikingly aligned with the original track, further reinforcing the presence of multimodal memorization. Similarly, for Jingle Bells Veo 3 consistently generated music that was melodically and rhythmically faithful to the original. This highlights the model’s strong tendency to regurgitate memorized content when exposed to exact training examples, extending lyric-based memorization across both audio and visual outputs.
4.5 Ablation Study
@@
To better understand the mechanisms underlying memorization in L2S and T2V models, we conduct a series of controlled ablation studies that isolate the effects of phonetic similarity versus verbatim content. By varying genre prompts, lyric fidelity, and phonetic perturbations across matched and mismatched inputs, we aim to disentangle the respective contributions of surface form, semantic content, and phonetic structure in triggering memorized generations. These studies expose the robustness and modality-transferability of memorization behaviors in modern generative models.
Genre Prompt Variation.
Even without any stylistic conditioning, YuE reproduces audio that closely aligns with training data when the lyrics match known examples. For instance, in the Mandarin song 天后 (Andrew Tan), AudioJudge assigns strong scores (melody = 0.88, rhythm = 0.85), while MiRA reports CLAP = 0.638 and CoverID = 0.300, indicating overlap with memorized content. This pattern mirrors MiRA’s earlier observations of lyric-driven leakage, with AudioJudge now confirming that acoustic structure is also faithfully preserved under verbatim prompting. In addition, supplying the correct genre tag amplifies memorization. For example, 光辉岁月 (Beyond) retains high melodic and rhythmic fidelity (0.95 / 0.90), with MiRA reporting CLAP = 0.731 and CoverID = 0.401. Likewise, 海阔天空 (Beyond) achieves nearly identical scores (melody = 0.95, rhythm = 0.92, CLAP = 0.767), showing that genre alignment neither reduces nor meaningfully alters memorized outputs when lyrics remain unchanged (Table 7 of Appendix C.2).
Same Song, Different Genre.
To test whether YuE responds more strongly to stylistic prompts or lyric memorization, we generated 后来 (Rene Liu) under four genre conditions (Table 8). Despite prompts ranging from inspiring pop to gentle piano ballad, AudioJudge and MiRA scores remain tightly clustered (melody = 0.90–0.95, rhythm = 0.75–0.92, CLAP = 0.785–0.858, CoverID = 0.291–0.570). These stable results indicate that genre conditioning has limited influence over musical structure, with YuE’s generations overwhelmingly anchored to the lyrics themselves, further suggesting a strong lyric-driven overfitting to training data (Table 8 of Appendix C.2).
5 Discussion
Table 5: Cosine similarity between original and modified lyrics across YuE/GPT embeddings.
Song Embedding Cosine Sim.
Lose Yourself YuE 0.976
GPT 0.746
DNA YuE 0.960
GPT 0.725
APT YuE 0.513
GPT 0.755
Why do phoneme-preserving prompts trigger such strong memorization across both audio and video generation models? We hypothesize that this phenomenon arises not merely from overfitting to training data, but from the central role that lyrics and rhythm play in the structure of the songs we evaluated. In particular, the rap and iconic pop we tested are characterized by tightly coupled lyrical phrasing, rhyme schemes, and rhythmic repetition. In these genres, the lyrics are not peripheral embellishments but serve as a core driver of musical identity. When this structure is mimicked, even through semantically nonsensical phrases, models may still activate memorized patterns tied to rhythm, syllabic stress, or acoustic cadence.
This interpretation is supported by embedding analyses (Table 5), which show that phoneme-preserving variants of rap tracks remain highly similar in YuE embeddings, reflecting the model’s reliance on rhythmic–phonetic alignment over semantic content. In contrast, genres where melody is the primary driver, such as modern K-Pop, showed lower sensitivity: phoneme-based prompts did not reproduce memorized outputs, despite the models’ strong performance in generating these songs.
These findings suggest that memorization in multi-modal generative systems is not merely a function of lexical overlap, but rather depends on the alignment between phonetic rhythm and musical phrasing. This adds a new dimension to the risk landscape for L2S models: even inputs that look safe at the text level may activate memorized content when they implicitly match the rhythmic fingerprint of songs seen during training. As generative systems scale, future defenses must consider not only token-level similarity, but also latent rhythmic and phonetic structure as potential leakage channels.
6 Conclusion
In this work, we introduce Adversarial PhoneTic Prompting (APT) attack, which exposes a new memorization vulnerability in L2S and T2V generation models. By altering lyrics to preserve phonetic structure while discarding semantics, we show that models like SUNO, YuE, and Veo 3 can reproduce memorized musical and visual content with high fidelity. These results highlight the model’s sensitivity to sub-lexical rhythm and cadence, revealing that phonetic cues alone—particularly in rhythmically structured genres like rap and iconic pop music—can serve as implicit triggers for memorization without lexical overlap or explicit cues. These findings expose an emerging risk in text-to-audio and transcript-conditioned generation pipelines, where phonetic form acts as a latent key to stored content. The success of our attack suggests that the demonstrated memorization behavior may emerge in transcript-conditioned generative systems, and we leave further investigation in this space of multi-modal generation for future work. As these models continue to be deployed in commercial and creative workflows, our results underscore the urgent need for new evaluation and safety frameworks that account for phonetic, rhythmic, and multimodal leakage paths, not just semantic or token-based similarity.
Appendix
Appendix A AudioJudge Heatmap
Figure 3: AudioJudge Similarity Heatmaps. We evaluate pairwise melody and rhythm similarity between original and generated songs using AudioJudge across four categories: (1) Mandarin, (2) Cantonese, and (3) other English songs. Each heatmap cell shows the overall similarity score (0–100) between an original and generated song. Green indicates high similarity (80–100), yellow moderate (40–79), and red low similarity (0–39). Diagonal cells reflect self-pairing scores (i.e., original with phoneme-modified versions of the same song). The distribution of scores confirms that AudioJudge does not assign uniformly high scores across all comparisons, but rather discriminates meaningfully based on melodic and rhythmic correspondence. This supports its reliability as an evaluative tool for music generation similarity.
Figure 3 provides a comprehensive visualization of AudioJudge’s similarity assessments across a diverse set of original and generated songs in four different evaluation scenarios: phoneme-modified English (top-left), Mandarin (top-right), English (bottom-left), and Cantonese (bottom-right). Each heatmap cell reflects the overall similarity score (range: 0–100), which aggregates melody and rhythm similarity scores produced by the GPT-4o-audio model given the AudioJudge prompt. To interpret the heatmaps: (i) Green cells (80–100) represent high similarity, (ii) Yellow cells (40–79) moderate similarity, and (iii) Red cells (0–39) low similarity.
Appendix B APT & AVT attack on Veo 3 Generation Example
(a) "Lose Yourself" Original MV Video Frame
(b) "Lose Yourself" Veo 3 Generation
(c) "Jingle Bell" Veo 3 Generation
Figure 4: Comparison between Veo 3-generated visuals through APT and AVT attacks.
Appendix C Additional Experimental Results
C.1 APT Attack: Christmas Songs Results